OpenStudy (anonymous):

simplify the expression, if possible.... log 0.1

OpenStudy (anonymous):

10 to the power of what has the result of 0.1?

OpenStudy (raden):

log 0.1 = log (1/10) = log 1 - log 10 = 0-1 = ......

OpenStudy (anonymous):

hm.. let's think of this way log itself is base of 10. and log x = 10^x so you can find the answer from now

OpenStudy (anonymous):

I have no clue lol...I'm going to review my study plan really quick

OpenStudy (anonymous):

correction: log x in exponential form = 10^x